### v2

@@ Abstract
The authors have no relevant competing interests to disclose. No funds, grants, or other support was received.
@@ Acknowledgements
The datasets generated during and/or analyzed during the current study are available in the FigShare repository, https://doi.org/10.6084/M9.FIGSHARE.22591513.
@@ 1 Introduction
The concept of sustainability has gained notable importance not only in general societal debates, but also regarding the activities of businesses. When considering the impact of business and their operation and processes, traditionally, the focus is put on performance in a financial or operational sense [31]. However, the scientific consensus about the relation between human activity and the existence and acceleration of human-made climate change and alterations of ecosystems underlines the necessity of considering environmental impact. The UN’s Sustainable Development Goals, released as a “blueprint to achieve a better and more sustainable future for all” [77], state that a decoupling of environmental degradation and resource use from economic growth and associated production and consumption patterns is required. According to the report, unsustainable human activities have significantly degraded the Earth’s ecosystems on a global scale, thereby endangering the ecological foundations of society. Toxic substances introduced into the environment as a result of e.g. wasteful manufacturing, play a significant role in the endangerment of biodiversity and promotion of risks to human health — a loss of biodiversity is closely linked to threats towards human life and its surrounding environment [67]. Further, industrial activity and production of heat and electricity, have been identified as key contributors to climate change [40]. Moreover, current goals for conserving and sustainably using nature, and achieving global sustainability, cannot be met by current trajectories, requiring transformative changes across economic, social, political, and technological factors [11].
@@
Concretely, SOPA contributes methods for assessing the impact of business processes on environmental sustainability by assigning impact measures based on LCA to activities with ABC, and proposes concrete measures for sustainability-oriented process re-design via business process simulation. In contrast to existing approaches, SOPA aims at a holistic consideration of the environmental impact through LCA, instead of focusing on a few explicit impact dimensions, and more explicitly covers mechanisms for re-design. In this work, we follow a design science research methodology [88, 59] in the design and evaluation of the artifact, SOPA. The evaluation is based on business process simulation [60] and a prototypical implementation, and applied to a real-world case study. We show how SOPA is able to be leveraged for process re-design by evaluating various re-design scenarios for their environmental impact and determining the least impactful variant.
The following section presents background information on sustainability, ABC, LCA, BPM and business process simulation. Then, the methodology of design science, used in this work, is detailed (Section 3). Subsequently, the resulting framework SOPA is presented and evaluated (Section 4). Finally, the results, as well as practical and theoretical considerations for BPM, are discussed (Section 5).
@@ 2.1 Sustainability
The term of sustainability has been established in discussions about viable long-term development, be it economic, social, or ecological. In 1987, the Brundtland Report of the World Commission on Environment and Development from the United Nations formulated sustainability as ensuring that the needs of the present are met without compromising the ability of future generations to meet their own needs [6]. Since then, some estimated three-hundred definitions of sustainability and sustainable development have emerged in environmental management and other disciplines [42]. This underlines that, while being a term of everyday use, defining the concept itself is complex, and a general understanding and uniform definition is still missing [62]. Some authors, such as [42] argue, therefore, that an unambiguous definition of sustainability and an authoritative framework that can be used to assess policies and practice w.r.t. sustainability is needed.
@@ 2.3.1 Business Process Simulation
To facilitate this simulation and analysis, a business process simulator is required [60]. This tool aims at simulating the execution of process instances based on input provided by a process analyst (e.g., resource information, branch probability, number of processes instances to simulate, and other parameters [51]), which is usually done in the form of a simulation configuration. The result of that simulation is artificial information in the form of, e.g., an event log and a performance report about the process, which provides, among others, information about probable throughput time, process costs and resource utilization of the current process design [12, 89]. The resulting information can then be used to iteratively drive the re-design of the considered process. However, it should be noted that BPS abstracts from the real-world behavior of the considered process, and, due to its stochastic and qualitative nature, depends greatly on the input parameters. Thus, they should be chosen with care and, if need be, validated externally [12]. An alternative approach is the use of historic data from logs [69], which can however only deliver reliable results if the process version from which the data stems is similar enough to the version that is simulated. In practice, some frameworks for process re-design explicitly relying on BPS have been outlined or implemented (e.g., [50]) – however, none of them include a notion of sustainability-oriented process re-design, which is a shortcoming that SOPA aims to address.
@@ 2.3.2 Activity-based Costing
ABC contains several key concepts, which are illustrated in Fig. 4. Cost objects represent the overall cost of the system under investigation, such as a product or a process. Activities that contribute to that product or process consume resources, which, through resource drivers, incur resource costs. In the context of BPM, resources are usually understood to be, among others, human resources, machines, vehicles, or materials, which greatly contribute to the execution of individual activities and process instances [39] — this is a perspective similar to that in ABC [28]. Resource drivers represent the demand placed on resources by an activity [46], and could be, e.g., the time an employee has spent on a specific activity, depreciation of the equipment involved in that activity, the supplies used, etc. Thus, by considering what resources contribute to an activity, and how the activity places demands on them, the resource cost of that activity can be determined. Cost drivers determine the cost of the operational activities to which they are linked. They be, e.g., the number of products ordered, number of parts required, etc. [46]. By considering cost drivers of activities and resource costs imposed on resources by an activity, the activity costs can be determined. Activity drivers represent the demand placed on activities, e.g., the number of times an activity needs to be executed, or the duration of it. With these activity drivers, the actual activity cost can be allocated to the overall cost object [28, 46]. Activity drivers could be, for example, the number of deliveries made to a customer, thus determining the overall cost for the customer in question [28]. When all activity costs have been considered and allocated to the cost object, the overall cost of the system has been determined.
@@ 3.2 System Development
The system functionalities and requirements to be developed, are implicitly part of the SOPA framework designed with DSR (see Section 3.1), which the system aims to implement on a technical level. Based on these requirements, a concrete system architecture with specific components that have specific functionalities is outlined in Section 4.2. This includes components such as a concrete BPS engine, and a tool for analyzing and calculating impact measures. Further, the process used to carry out the respective system functions (i.e., to conduct sustainability-oriented process analysis and re-design), is illustrated in Section 4.2. Finally, the system is built and evaluated with software components developed specifically for this purpose in Section 4.3. These resulting components are available as open-source software111See https://anonymous.4open.science/r/sustainability-analysis-tool-E0DE and https://anonymous.4open.science/r/scylla-3192 [Accessed: July 17, 2024, to be replaced with a non-blind link after review]. Thus, the evaluation of the concrete system also serves as a general evaluation of SOPA, following the penultimate step of DSR, by demonstrating its applicability and utility in a practical setting.
@@ 3.3 Case Study — Data Collection
The LCA analysis was conducted with mobius333See https://mobius.ecochain.com/ [Accessed: July 17, 2024] a web-based tool for LCA analyses. The concrete LCA method used was EF 3.0 (Ecoinvent v 3.8 Cut-Off) [18], with normalization and weighting (see Section 2.2). Multiple impact scores were calculated in reliance on the Ecoinvent database v3.8444See https://ecoinvent.org/the-ecoinvent-database/data-releases/ecoinvent-3-8/ [Accessed: July. 17, 2024]. The data elicited for this case study is available online555See https://figshare.com/s/b0837223b9bf5e109859 [Accessed: July 17, 2024, to be replaced with a permanent non-blind DOI after review]. Figure 7 displays that process, which will be used subsequently as a running example to explain the application of SOPA. A higher resolution of the process model is available online at the same location.
@@ 4.1 Framework and Metamodel
SOPA proposes a three-step approach for analyzing business processes for their environmental impact and their sustainability-oriented re-design. When applying SOPA to analyze business processes for their environmental impact and to re-design them accordingly, two possible scenarios exist. Either an existing event log, which contains execution data of real-world process instances, is analyzed and used for reasoning about improvement potentials (log-based analysis), or a process model is, through utilization of BPS, assessed (model-based analysis). This leads to the following three central steps of SOPA:
@@ 4.2 System Development
The system’s architecture consists of multiple components used in concert by a process analyst to elicit environmental impact information, consisting of average process instance and activity environmental costs, of a business process under investigation. Figure 12 illustrates the different components that are involved. First, the process analyst can provide a process model annotated with abstract environmental cost driver information and a simulation configuration with cost variants to a fork of the process simulation engine Scylla777See https://anonymous.4open.science/r/scylla-3192 [Accessed: July 17, 2024, to be replaced with a non-blind link after review] (see [60]), that has been modified to accommodate the SOPA-specific concepts. The resulting (simulated) event log is used by the Analysis Tool888See https://anonymous.4open.science/r/sustainability-analysis-tool-E0DE [Accessed: July 17, 2024, to be replaced with a non-blind link after review], together with the simulation configuration (containing cost variants and the corresponding concrete environmental cost drivers) and process model provided to Scylla, to calculate the relevant impact information. Alternatively, an event log containing the relevant information that has been extracted out of a real-world information system could be used as well. The impact information can then be used by the process analyst to either put the hypothesized changes into practice, or to drive a new re-design iteration.
@@ 5.1.2 SOPA
First, it should be noted that the quality of the results elicited with SOPA depends on the completeness of the input, be it a simulated event log or an extracted one — the more complete and close-to-real-world the captured behavior and the impact information is, the more accurate the assessment with SOPA is. To increase the exactness of simulation, stochastic information about the process can either be estimated by process experts or identified based on the analysis of process execution logs [89] to increase accuracy. The more in-depth an analyst models the process (e.g., across multiple participants), the more the impact can be attributed to different participants, attributing concrete responsibilities.
@@ 5.2.2 Sustainability through BPM?
Following [30], the governance of nature would require, instead of the prevalent monetized cost-benefit analyses, multidimensional evidence. Unlocking path dependencies towards sustainability arguably is hindered by inequalities inherent to market logic, laws, and institutions, that are being created, overlooked, or disguised through undifferentiated cost–benefit and growth analyses [30]. Further, sustainability solutions may not lie in increasing the efficiency of technologies and the effectiveness of economic incentives, but instead may need to be facilitated by organizational setups and sociocultural frameworks defining the purpose of technologies and economic instruments [30]. Thus, while SOPA is a contribution within BPM to enable further considerations of sustainability with it, the overall role of BPM itself might need to be investigated closely for how it has internalized mechanisms that might hinder adequately addressing unsustainable path dependencies, and how it might be leveraged for sustainability solutions within institutions.
@@ Appendix A Illustrating Example
In the following, we provide further examples of the concepts we formalized in Section 4.1. We stress that these examples primarily serve to illustrate the concepts, and are explicitly not meant for completeness – for this, we refer to the case study in Section 4.3 and the corresponding supplementary material available online.111111https://figshare.com/s/b0837223b9bf5e109859 [Accessed: July 17, 2024, to be replaced with a permanent non-blind DOI after review]

### v3

@@ Abstract
The authors have no relevant conflict of interest to disclose. Open access funding was provided by the Technical University of Munich. No further funds, grants, or other support was received.
@@ Acknowledgements
The datasets generated during and/or analyzed during the present study are available in a FigShare repository at https://doi.org/10.6084/M9.FIGSHARE.22591513.
@@ 1 Introduction
The concept of sustainability has gained notable importance not only in general societal debates, but also regarding the activities of businesses. When considering the impact of business and their operation and processes, traditionally, the focus is put on performance in a financial or operational sense [31]. However, the scientific consensus about the relation between human activity and the existence and acceleration of human-made climate change and alterations of ecosystems underlines the necessity of considering environmental impact. The UN’s Sustainable Development Goals, released as a “blueprint to achieve a better and more sustainable future for all” [77], state that a decoupling of environmental degradation and resource usage from economic growth and associated production and consumption patterns is required. According to the report, unsustainable human activities have significantly degraded the Earth’s ecosystems on a global scale, thereby endangering the ecological foundations of society. Toxic substances introduced into the environment as a result of e.g. wasteful manufacturing, play a significant role in the endangerment of biodiversity and promotion of risks to human health — a loss of biodiversity is closely linked to threats towards human life and its surrounding environment [67]. Further, industrial activity and production of heat and electricity have been identified as key contributors to climate change [40]. Moreover, “current goals for conserving and sustainably using nature, and achieving global sustainability, cannot be met by current trajectories”, requiring “transformative changes across economic, social, political, and technological factors” [11].
@@
Concretely, SOPA contributes methods for assessing the impact of business processes on environmental sustainability by assigning impact measures based on LCA to activities with ABC, and proposes concrete measures for sustainability-oriented process re-design via business process simulation. In contrast to existing approaches, SOPA aims at a holistic consideration of the environmental impact through LCA, instead of focusing on a few explicit impact dimensions, and more explicitly covers mechanisms for re-design. In this work, we follow a design science research methodology [88, 59] in the design and evaluation of the artifact, SOPA. The evaluation is based on business process simulation [60] and a prototypical implementation, and is applied to a real-world case study. We show how SOPA is able to be leveraged for process re-design by evaluating various re-design scenarios for their environmental impact and determining the least impactful variant.
The following section presents background information on sustainability, ABC, LCA, BPM and business process simulation. Then, the methodology of design science, used in this work, is detailed (Section 3). Subsequently, the resulting framework of SOPA is presented and evaluated (Section 4). Finally, the results, as well as practical and theoretical considerations for BPM, are discussed (Section 5).
@@ 2.1 Sustainability
The term of sustainability has been established in discussions about viable long-term development, be it economic, social, or ecological. In 1987, the Brundtland Report of the World Commission on Environment and Development from the United Nations formulated sustainability as ensuring that the needs of the present are met without compromising the ability of future generations to meet their own needs [6]. Since then, some estimated three-hundred definitions of sustainability and sustainable development have emerged in environmental management and other disciplines [42]. This underlines that, while being a term of everyday use, defining the concept itself is complex, and a general understanding and uniform definition is still missing [62]. Some authors, such as [42] argue, therefore, that an unambiguous definition of sustainability and an authoritative framework that can be used to assess policies and practice w.r.t. sustainability are needed.
@@ 2.3.1 Business Process Simulation
To facilitate this simulation and analysis, a business process simulator is required [60]. This tool aims at simulating the execution of process instances based on input provided by a process analyst (e.g., resource information, branch probability, number of processes instances to simulate, and other parameters [51]), which is usually done in the form of a simulation configuration. The result of that simulation is artificial information in the form of, e.g., an event log and a performance report about the process, which provides, among others, information about probable throughput time, process costs and resource usage of the current process design [12, 89]. The resulting information can then be used to iteratively drive the re-design of the considered process. However, it should be noted that BPS abstracts from the real-world behavior of the considered process, and, due to its stochastic and qualitative nature, depends greatly on the input parameters. Thus, they should be chosen with care and, if need be, validated externally [12]. An alternative approach is the use of historic data from logs [69], which can however only deliver reliable results if the process version from which the data stems is similar enough to the version that is simulated. In practice, some frameworks for process re-design explicitly relying on BPS have been outlined or implemented (e.g., [50]) – however, none of them include a notion of sustainability-oriented process re-design, which is a shortcoming that SOPA aims to address.
@@ 2.3.2 Activity-based Costing
ABC contains several key concepts, which are illustrated in Fig. 4. Cost objects represent the overall cost of the system under investigation, such as a product or a process. Activities that contribute to that product or process consume resources, which, through resource drivers, incur resource costs. In the context of BPM, resources are usually understood to be, among others, human resources, machines, vehicles, or materials, which greatly contribute to the execution of individual activities and process instances [39] — this is a perspective similar to that in ABC [28]. Resource drivers represent the demand placed on resources by an activity [46], and could be, e.g., the time an employee has spent on a specific activity, depreciation of the equipment involved in that activity, the supplies used, etc. Thus, by considering what resources contribute to an activity, and how the activity places demands on them, the resource cost of that activity can be determined. Cost drivers determine the cost of the operational activities to which they are linked. They could be, e.g., the number of products ordered, number of parts required, etc. [46]. By considering cost drivers of activities and resource costs imposed on resources by an activity, the activity costs can be determined. Activity drivers represent the demand placed on activities, e.g., the number of times an activity needs to be executed, or the duration of it. With these activity drivers, the actual activity cost can be allocated to the overall cost object [28, 46]. Activity drivers could be, for example, the number of deliveries made to a customer, thus determining the overall cost for the customer in question [28]. When all activity costs have been considered and allocated to the cost object, the overall cost of the system has been determined.
@@ 3.2 System Development
The system functionalities and requirements to be developed, are implicitly part of the SOPA framework designed with DSR (see Section 3.1), which the system aims to implement on a technical level. Based on these requirements, a concrete system architecture with specific components that have specific functionalities is outlined in Section 4.2. This includes components such as a concrete BPS engine, and a tool for analyzing and calculating impact measures. Further, the process used to carry out the respective system functions (i.e., to conduct sustainability-oriented process analysis and re-design), is illustrated in Section 4.2. Finally, the system is built and evaluated with software components developed specifically for this purpose in Section 4.3. These resulting components are available as open-source software111See https://github.com/INSM-TUM/sustainability-analysis-tool and https://github.com/INSM-TUM/sustainability-scylla-extension [Accessed: July 17, 2024]. Thus, the evaluation of the concrete system also serves as a general evaluation of SOPA, following the penultimate step of DSR, by demonstrating its applicability and utility in a practical setting.
@@ 3.3 Case Study — Data Collection
The LCA analysis was conducted with mobius333See https://mobius.ecochain.com/ [Accessed: July 17, 2024] a web-based tool for LCA analyses. The concrete LCA method used was EF 3.0 (Ecoinvent v 3.8 Cut-Off) [18], with normalization and weighting (see Section 2.2). Multiple impact scores were calculated in reliance on the Ecoinvent database v3.8444See https://ecoinvent.org/the-ecoinvent-database/data-releases/ecoinvent-3-8/ [Accessed: July. 17, 2024]. The data elicited for this case study is available online555See https://figshare.com/s/b0837223b9bf5e109859 [Accessed: July 17, 2024]. Figure 7 displays that process, which will be used subsequently as a running example to explain the application of SOPA. A higher resolution of the process model is available online at the same location.
@@ 4.1 Framework and Metamodel
SOPA proposes a three-step approach for analyzing business processes for their environmental impact and their sustainability-oriented re-design. When applying SOPA to analyze business processes for their environmental impact and to re-design them accordingly, two possible scenarios exist. Either an existing event log, which contains execution data of real-world process instances, is analyzed and used for reasoning about improvement potentials (log-based analysis), or a process model is, through usage of BPS, assessed (model-based analysis). This leads to the following three central steps of SOPA:
@@ 4.2 System Development
The system’s architecture consists of multiple components used in concert by a process analyst to elicit environmental impact information, consisting of average process instance and activity environmental costs, of a business process under investigation. Figure 12 illustrates the different components that are involved. First, the process analyst can provide a process model annotated with abstract environmental cost driver information and a simulation configuration with cost variants to a fork of the process simulation engine Scylla777See https://github.com/INSM-TUM/sustainability-scylla-extension [Accessed: July 17, 2024] (see [60]), that has been modified to accommodate the SOPA-specific concepts. The resulting (simulated) event log is used by the Analysis Tool888See https://github.com/INSM-TUM/sustainability-analysis-tool [Accessed: July 17, 2024], together with the simulation configuration (containing cost variants and the corresponding concrete environmental cost drivers) and process model provided to Scylla, to calculate the relevant impact information. Alternatively, an event log containing the relevant information that has been extracted out of a real-world information system could be used as well. The impact information can then be used by the process analyst to either put the hypothesized changes into practice, or to drive a new re-design iteration.
@@ 5.1.2 SOPA
First, it should be noted that the quality of the results elicited with SOPA depends on the completeness of the input, be it a simulated event log or an extracted one — the more complete and close-to-real-world the captured behavior and the impact information is, the more accurate the assessment with SOPA is. To increase the exactness of simulation, stochastic information about the investigated process can either be estimated by process experts or identified based on the analysis of event logs [60, 89] to increase accuracy. The more in-depth an analyst models the process (e.g., across multiple participants), the more the impact can be attributed to different participants, attributing concrete responsibilities.
@@ 5.2.2 Sustainability through BPM?
Following [30], the governance of nature would require, instead of the prevalent “monetized cost-benefit analyses”, multidimensional evidence. Unlocking path dependencies towards sustainability arguably is hindered by inequalities inherent to “market logic, laws, and institutions [that are being] created, overlooked, or disguised [through] undifferentiated cost–benefit and growth analyses” [30]. Further, sustainability solutions may not lie in increasing the efficiency of technologies and the effectiveness of economic incentives, but instead may need to be facilitated by organizational setups and sociocultural frameworks defining the purpose of technologies and economic instruments [30]. Thus, while SOPA is a contribution within BPM to enable further considerations of sustainability with it, the overall role of BPM itself might need to be investigated closely for how it has internalized mechanisms that might hinder adequately addressing unsustainable path dependencies, and how it might be leveraged for sustainability solutions within institutions.
@@ Appendix A Illustrating Example
In the following, we provide further examples of the concepts we formalized in Section 4.1. We stress that these examples primarily serve to illustrate the concepts, and are explicitly not meant for completeness – for this, we refer to the case study in Section 4.3 and the corresponding supplementary material available online.111111https://doi.org/10.6084/M9.FIGSHARE.22591513 [Accessed: July 17, 2024]